# Commute Graph Neural Networks

### v1

@@ C.2 Data Statistics
C.3 Hyperparameter Settings
For our model, we tune the hyperparameters based on the highest average validation accuracy. We utilize the randomized truncated SVD algorithm for computing the Moore-Penrose pseudoinverse of matrix $\mathcal{R}$, setting the required rank $q$ to 5 for all datasets. The learning rate $lr$ is selected from $\{0.01,0.005\}$, and the weight decay $wd$ from $\{0,5e-5,5e-4\}$. In the model architecture, the number of layers $L$ vary among $\{1,2,3,4,5\}$ and the dimension $d^{\prime}$ is selected from $\{32,64,128,256,512\}$. The comprehensive hyperparameter configurations for CGNN are detailed in Table 5.
Table 5: Hyperparameters specifications.
Dataset $lr$ $wd$ $L$ $d^{\prime}$
Squirrel 0.005 0 5 128
Chameleon 0.01 0 4 128
CoraML 0.01 0 2 64
Citeseer 0.01 0 2 128
AM-Photo 0.005 0 2 512
Appendix D Additional Experiments
Table 6: Link prediction results.
@@ Appendix D Additional Experiments
Direction prediction Existence prediction Method Citeseer CoraML Citeseer CoraML GCN 77.69$\pm$0.73 75.68$\pm$0.97 70.27$\pm$2.12 80.67$\pm$1.08 GAT 76.09$\pm$0.09 50.02$\pm$0.05 50.82$\pm$0.36 50.00$\pm$0.01 GraphSAGE 75.16$\pm$0.91 70.07$\pm$1.43 70.59$\pm$1.74 69.32$\pm$0.58 APPNP 79.23$\pm$0.35 78.64$\pm$0.70 81.14$\pm$1.21 85.46$\pm$0.80 MixHop 79.40$\pm$0.42 78.62$\pm$1.25 80.11$\pm$1.93 85.21$\pm$0.90 GPRGNN 79.15$\pm$0.35 77.88$\pm$1.30 78.15$\pm$0.24 83.83$\pm$0.74 GCNII 78.60$\pm$0.69 75.51$\pm$1.90 79.48$\pm$1.74 75.51$\pm$1.90 DGCN 82.38$\pm$0.55 74.13$\pm$1.07 81.11$\pm$2.28 80.30$\pm$1.25 DiGCN 83.06$\pm$0.70 74.92$\pm$0.88 85.08$\pm$1.26 81.06$\pm$1.16 MagNet 79.36$\pm$0.44 77.71$\pm$1.04 88.20$\pm$1.05 87.21$\pm$0.72 DiGCL 81.89$\pm$0.76 76.05$\pm$0.97 82.96$\pm$1.32 84.23$\pm$1.26 DirGNN 82.09$\pm$0.88 77.93$\pm$0.76 85.18$\pm$1.02 83.37$\pm$1.83 CGNN 84.97$\pm$0.71 80.33$\pm$0.90 85.27$\pm$1.30 82.08$\pm$1.22
D.1 Link Prediction
For link prediction, we compare our method with all baselines on Citeseer and CoraML datasets. Our experimental setup follows the protocol established by (Zhang et al., 2021), considering two distinct types of link prediction tasks for performance evaluation. The first task involves predicting the direction of edges between pairs of nodes, while the second focuses on predicting the existence of edges. For both tasks, we partition the data by removing 15% of the edges for testing and 5% for validation, with the remaining edges used for training. The results, detailed in Table 6, reveal that CGNN excels in the direction prediction task. Meanwhile, in the edge existence prediction task, MagNet yields the best results, with CGNN also achieving comparable performance. These findings show CGNN’s effectiveness in accurately capturing directed structural patterns in graphs.
D.2 Sensitivity analysis
We investigate the sensitivity of CGNN to key hyperparameters that influence its performance, specifically focusing on the number of layers $L$ and the dimension of the hidden layer $d^{\prime}$. We explore a range of values for $L$, considering $\{1,2,3,4,5\}$, and for $d^{\prime}$, considering $\{32,64,128,256,512\}$. From Fig. 5, we observe that we observe that CGNN achieves optimal performance with $L=5$ and $d^{\prime}=128$ on Squirrel, and with $L=2$ and $d^{\prime}=64$ on CoraML. This suggests that deeper models are necessary to effectively aggregate valuable information in heterophilic graphs, whereas in homophilic graphs, leveraging local neighborhood information is generally adequate.
Figure 5: Sensivisity analysis on Squirrel and CoraML.

### v2

@@ 6.2 Data Statistics
Snap-Patents 2,923,922 13,975,791 269 5 0.22
Roman-Empire 22,662 44,363 300 18 0.05
Arxiv-Year 169,343 1,166,243 128 40 0.22
Table 2: Node classification results. Accuracy (%) with standard deviation for 10 runs. We highlight/underline the best/second-best method. For general GNN and non-local GNN baselines, we conduct experiments on both symmetrized versions and their directed counterparts, reporting better results from these two settings. OOM indicates out-of-memory.
@@
Method Squirrel Chameleon Citeseer CoraML AM-Photo Snap-Patents Roman-Empire Arxiv-Year GCN 52.43$\pm$2.01 67.96$\pm$1.82 66.03$\pm$1.88 70.92$\pm$0.39 88.52$\pm$0.47 51.02$\pm$0.06 73.69$\pm$0.74 46.02$\pm$0.26 GAT 40.72$\pm$1.55 60.69$\pm$1.95 65.58$\pm$1.39 72.22$\pm$0.57 88.36$\pm$1.25 OOM 49.18$\pm$1.35 45.30$\pm$0.23 GraphSAGE 41.61$\pm$0.74 62.01$\pm$1.06 66.81$\pm$1.38 74.16$\pm$1.55 89.71$\pm$0.57 67.45$\pm$0.53 86.37$\pm$0.80 55.43$\pm$0.75 APPNP 51.91$\pm$0.56 45.37$\pm$1.62 66.90$\pm$1.82 70.31$\pm$0.67 87.43$\pm$0.98 51.23$\pm$0.54 72.96$\pm$0.38 50.31$\pm$0.42 MixHop 43.80$\pm$1.48 60.50$\pm$2.53 56.09$\pm$2.08 65.89$\pm$1.50 87.17$\pm$1.34 41.22$\pm$0.19 50.76$\pm$0.14 45.30$\pm$0.26 GPRGNN 50.56$\pm$1.51 66.31$\pm$2.05 61.74$\pm$1.87 73.31$\pm$1.37 90.23$\pm$0.34 40.19$\pm$0.03 64.85$\pm$0.27 45.07$\pm$0.21 GCNII 38.47$\pm$1.58 63.86$\pm$3.04 58.32$\pm$1.93 64.84$\pm$0.71 83.40$\pm$0.79 48.09$\pm$0.09 74.27$\pm$0.13 57.36$\pm$0.17 DGCN 37.16$\pm$1.72 50.77$\pm$3.31 66.37$\pm$1.93 75.02$\pm$0.50 87.74$\pm$1.02 OOM 51.92$\pm$0.43 OOM DiGCN 33.44$\pm$2.07 50.37$\pm$4.31 64.99$\pm$1.72 77.03$\pm$0.70 88.66$\pm$0.51 OOM 52.71$\pm$0.32 48.37$\pm$0.19 MagNet 39.01$\pm$1.93 58.22$\pm$2.87 65.04$\pm$0.47 76.32$\pm$0.10 86.80$\pm$0.65 OOM 88.07$\pm$0.27 60.29$\pm$0.27 DUPLEX 57.60$\pm$0.98 61.25$\pm$0.94 67.60$\pm$0.72 72.26$\pm$0.71 87.80$\pm$0.82 66.54$\pm$0.11 79.02$\pm$0.08 64.37$\pm$0.27 DiGCL 35.82$\pm$1.73 56.45$\pm$2.77 67.42$\pm$0.14 77.53$\pm$0.14 89.41$\pm$0.11 70.65$\pm$0.07 87.94$\pm$0.10 63.10$\pm$0.06 DirGNN 75.19$\pm$1.26 79.11$\pm$2.28 66.57$\pm$0.74 75.33$\pm$0.32 88.09$\pm$0.46 73.95$\pm$0.05 91.23$\pm$0.32 64.08$\pm$0.26 CGNN 77.61$\pm$1.87 79.54$\pm$1.82 70.27$\pm$0.66 77.06$\pm$0.67 90.41$\pm$0.32 72.33$\pm$0.16 92.89$\pm$0.51 66.97$\pm$0.15
6.3 Overall Results and Analysis
Table 2 reports the node classification results across eight digraph datasets. Our method CGNN achieves new state-of-the-art results on 6 out of 8 datasets, and comparable results on the remaining ones, validating the superiority of CGNN. We provide more observations as follows. Firstly, while non-local GNNs have the potential to cover the commute paths between adjacent nodes by stacking multiple layers, they do not consistently outperform general, shallow GNN models. It suggests that coarsely aggregating all nodes in commute paths is ineffective. The reason is that deeper models may aggregate excessive irrelevant information for the central node. Our goal is to encode mutual relationships between adjacent nodes by considering their commute times. Aggregating all nodes along the entire path introduces excessive information about other nodes unrelated to the direct relationship between the target nodes. Secondly, GNNs tailored for digraphs do not seem to bring substantial gains. Our results show that with careful hyper-parameter tuning, general GNNs can achieve results comparable to, or even better than, those tailored for digraphs, as evidenced in the Squirrel, Chameleon, and AM-Photo datasets. Thirdly, CGNN achieves state-of-the-art results on both homophilic and heterophilic digraph benchmarks. Notably, DirGNN also performs comparably on heterophilic graphs (e.g., Squirrel and Chameleon), supporting the findings of [7] that distinguishing edge directionality during message passing enables the central node to adaptively balance information flows from both heterophilic and homophilic neighbors, effectively mitigating the impact of heterophily. Moreover, CGNN, an enhanced version of DirGNN, further improves performance on these graphs by effectively incorporating commute times to refine the strength of relationships between nodes, enhancing model robustness under heterophily.
(a) Squirrel
(b) AM-Photo
Figure 3: Accuracy vs. running time on Squirrel and AM-Photo.
Application scope analysis.
Can commute times always enhance message passing on directed graphs? To answer this, we analyze the scope of use for CGNN based on Table 2. For example, on the Snap-Patents and CoraML dataset, we observed that adding commute time-based weights during message passing did not significantly enhance performance. Now we can analyze the reason from the perspective of dataset. CoraML is a directed citation network where nodes predominantly link to other nodes within the same research area. However, in such networks, reciprocal citations between two papers are impossible due to their chronological sequence. Consequently, mutual path dependencies do not exist, and thus, incorporating commute times to adjust neighbor weights might might (slightly) hurt performance. A similar situation exists with the Snap-Patents dataset, where each directed edge represents a citation from one patent to another, again indicating the absence of mutual path dependencies.
In conclusion, in networks like citation networks where mutual relationships inherently do not exist, applying commute times to enforce these relationships is unnecessary. Conversely, our model is particularly effective in networks like webpage networks and social networks—examples being Squirrel and AM-Photo—where mutual relationships are prevalent. For instance, in a social network, an ordinary user may follow a celebrity, creating a short path to the celebrity. However, the reverse path from the celebrity back to the user might be considerably longer. Therefore, in such networks, considering mutual relationships based on commute times can provide a more accurate description of node relationships.
Table 4: Accuracy and running time (s) of CGNN and CGNNppr.
Squirrel Chameleon Citeseer CoraML AM-Photo Method Acc. Time Acc. Time Acc. Time Acc. Time Acc. Time CGNN 77.61 102.93 79.54 112.24 70.27 92.26 77.02 127.15 90.41 121.80 CGNNppr 68.37 257.84 71.69 253.05 68.59 137.82 76.23 192.09 88.52 203.04
6.4 Efficiency Comparsion
6.5 Component Analysis
Table 5: Changes in commute times before and after rewiring.
CoraML Chameleon Roman-Empire $\delta$ 0.03280 0.00157 0.06242
Comparison between graph rewiring and PPR.
In Section 5.2, we construct a rewired graph $\widetilde{G}$ based on feature similarity to guarantee the irreducibility and aperiodicity. This approach introduces at most two additional edges per node, specifically targeting those with the highest feature similarity, while minimally altering the original graph structure to preserve semantic information. To investigate changes in commute times before and after rewiring, for the original graph, we use its largest connected component, removing absorbing nodes (i.e., nodes without outgoing edges) to ensure that we can compute meaningful and deterministic commute times. We denote the average normalized commute time for the original graph as $c_{orig}$; For the rewired graph, we directly compute the commute time and denote this average normalized value as $c_{rew}$. We then use $\delta=\frac{\|c_{orig}-c_{rew}\|_{2}}{\|c_{orig}\|}$ to quantify the changes, which can be intepretated as the proportion of commute information changed in the original graph. As shown in Table 5, the graph rewiring method can effectively preserve the original commute times of the graph.
In contrast, the classic PageRank transition matrix, defined as $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, achieves a similar objective but results in a completely connected graph $G_{pr}$. However, this approach tends to overlook the sparse structure of the original graph, which may alter the semantic information in the graph. Additionally, computing commute times using a dense transition matrix incurs a high computational cost. To validate the effectiveness of the rewiring approach over the PPR method, we conduct an experiment where $\widetilde{G}$ is replaced with $G_{pr}$ in the computation of commute-time-based proximity. We denote this variant as ‘CGNNppr’ and the results of accuracy and efficiency are reported in Table 4. The findings reveal that the PPR approach is suboptimal in terms of both accuracy and efficiency, thereby underscoring the effectiveness of our rewiring-based approach.
Table 6: Impact of directed structure.
Directed vs. Undirected.
To validate the critical role of directed structures in our model, we transform all directed edges into undirected ones by adding their reverse counterparts. This process results in a symmetric adjacency matrix, denoted as $\mathbf{A}_{\mathrm{sym}}$. Subsequently, the commute time is calculated based on the transition matrix derived from $\mathbf{A}_{\mathrm{sym}}$. We refer this variant as ‘CGNNsym’. Table 6 shows the accuracy of CGNN and CGNNsym on three datasets. We find that edge direction can significantly influence the prediction accuracy for our model.
7 Conclusion
Identifying and encoding asymmetric mutual path dependencies in directed graphs is essential for accurately representing real-world relationships between entities. In this work, we utilize the concept of commute time to assess the strength of relationships in directed graphs and introduce the Commute Graph Neural Network (CGNN) to incorporate node-wise commute time into node representations. To achieve this, we propose DiLap, a novel Laplacian formulation derived from the divergence of the gradient of signals on directed graphs, along with an efficient computational method for deterministic commute times. By integrating commute times into GNN message passing through neighbor weighting, CGNN effectively harnesses path asymmetry in directed graphs, thereby enhancing node representation learning. Our extensive experiments across eight directed graph datasets demonstrate that CGNN significantly outperforms existing methods.